Give the starting network address 181.173.40.0/24 subnet to make subnets that support a minimum 36 hosts each while making as many subnets as possible. What is the new mask in CIDR notation? Question 3 options:

a) 28

b) 30

c) 25

d) 27

e) 29

f) 26
Answer #1

Solution:

Given,

=>Network address = 181.173.40.0/24

=>Minimum number of hosts per subnet = 36

The answer will be an option,

(f) 26

Explanation:

Calculating minimum number of bits for host ID part:

=>Number of bits required for host ID part = ceil(log2(number of hosts))

=>Number of bits required for host ID part = ceil(log2(36))

=>Number of bits required for host ID part = ceil(5.16)

=>Number of bits required for host ID part = 6 bits

Calculating number of bits for subnetting:

=>As only 2 bits out of 8 bits are left for subnetting so 2 bits will be used for subnetting.

=>Subnet ID bits + net ID bits = 2 bits + 24 bits

=>Subnet ID bits + net ID bits = 26 bits

=>Subnet mask = 11111111.11111111.11111111.11000000/26

=>Subnet mask = 255.255.255.192/26

=>Hence in CIDR notation we will use /26"

=>Hence on the basis of above statements option (f) is correct and other options are incorrect because in other options minimum 36 hosts are not possible as 6 bits are required for host ID so maximum subnet ID and net ID part in total should be 26 bits and we are interested in maximum subnets so 26 bits is required.

I have explained each and every part with the help of statements attached to the answer above.